A sled weighing 60.0 N is pulled horizontally across snow so that the coefficient of kinetic friction between sled and snow is 0.100. A penguin weighing 70.0 N rides on the sled.If the coefficient of static fric- tion between penguin and sled is 0.700, find the maximum horizontal force that can be exerted on the sled before the penguin begins to slide off.
1
Expert's answer
2020-02-24T10:49:58-0500
"ma=F_{f,max}=\mu_smg\to a=\mu_s g"

"ma=F-F_f=F-\mu_kmg"

"F=(\mu_s+\mu_k)mg=(\mu_s+\mu_k)(W_s+W_p)"

"F=(0.7+0.1)(60+70)=104\ N"